l is the line x-2y+2=0.
m is the line 3x+y-8=0.
Use simultaneous equations to find the coordinates of P, the point of intersection of l and m.

Answer by Cromlix(4381)

Hi there,
x-2y+2=0.
sort out to
x - 2y = -2....(1)
......
3x+y-8=0.
sort out to
3x + y = 8......(2)
........
x - 2y = -2....(1)
3x + y = 8......(2)
Multiply (2) by 2
x - 2y = -2....(1)
6x + 2y = 16....(2)
Add (1) + (2)
7x = 14
x = 2
Substitute x = 2 into Eq(1)
x - 2y = -2
2 - 2y = -2
-2y = -2 - 2
-2y = -4
y = 2
Point of intersection of
l and m is (2,2)
Hope this helps:-)